# Generative Photomontage

### v2

@@ 1 Introduction
Our framework begins with a stack of images from ControlNet, generated by using the same text prompt and control inputs, and lets users choose parts they like from different images via simple brush strokes. Our key insight is that these images share common spatial structures from the same input condition, which can be leveraged for composition. We propose a novel technique that takes in the user’s brush strokes, segments the image parts in diffusion feature space, and then composites these parts during a final denoising process. Specifically, given users’ sparse scribbles, we formulate a multi-label graph-based optimization in diffusion feature space, grouping regions with similar diffusion features while satisfying user inputs. We then introduce a new feature injection and mixing method to composite the segmented regions. Our method accurately preserves the user-selected regions while harmoniously blending them together.
@@
We show visually compelling results on various applications and user workflows, including creating new appearance combinations, correcting shape misalignment, reducing artifacts, and improving prompt alignment. Our method outperforms existing pixel-space blending methods in preserving the realism and fidelity of local regions. Our code, data, and additional results are available on our webpage.
@@ 2 Related Work
Image editing.   Text-to-image diffusion models have enabled various editing tasks given reference images and text-instructions [51, 31, 34, 26, 80, 81, 87]. This is usually achieved through fine-tuning the model [40, 12] or modifying the denoising process of the diffusion model [51, 30, 55, 76, 14, 2, 35]. These methods often focus on the attention mechanism in the text-to-image model, which is crucial for determining the structure and text alignment of generated images [30, 57]. While we take inspiration from these works, our tasks and methods are different. Notably, MasaCtrl [14], Cross-Image Attention [2], and StyleAligned [31] focus on high-level style transfer, where local appearances are expected to change. We focus on blending a multi-image stack and preserving local appearances for greater user control. In contrast to their approaches, which perform full image feature injection, our method solves for and injects a composite of features from multiple images, computed from a graph-cut optimization. We show that our proposed technique performs better for this new task.
Controllable image generation.   Improving controllability and adherence to text instructions is critical for using these models as a collaborative tool. As a result, many recent works increase user control in the form of input conditions, such as sketch, depth map, bounding box, segmentation map, and reference image [84, 60, 5, 48, 42, 88, 56, 50, 27, 9]. Another line of work improves the existing text conditioning [17, 23, 7], constrains internal features [75, 6], or augments it through rich text editor [25]. However, these works aim to create a single correct image by directly constraining the output space of solutions. In contrast, we offer a complementary approach – we allow users to pick and choose exactly what they want from multiple generated images, giving them more fine-grained control without relying solely on the model to create a single perfect image in one shot.
@@ 3.2 Segmentation with Feature-Space Graph Cut
Discussion.   While it is possible to segment each image individually using off-the-shelf methods (e.g., SAM [43]), it often creates conflicts across the image stack where objects overlap, requiring extra user intervention to resolve conflicts. Instead, our multi-label graph cut optimization can automatically account for all images within the stack, assigning a unique label to each location.
@@ 4 Results
Figure 5: Appearance Mixing Results. Examples of using Generative Photomontage for creative exploration. (a, b) Users can combine different architectural elements (e.g., roof, windows, door) to form new architectural designs. (c) The user combines different colors of snakes to create a striking image. (d) The user selects the yellow beak from the second image and combines different vibrant colors in a bird’s feather to explore new looks. (e) The user combines different parts of a futuristic robot to form their favorite look.
Figure 6: Shape and Artifacts Correction Results. Our method can fix incorrect shapes and artifacts from ControlNet’s outputs, which often occur for uncommon input shapes. For example, (a) users can remove the extra rock at the Apple bite with a patch of grass from the second image, and (b) correct the shape and contour of the first waffle by selecting the background region of the second image. Users can also (c) correct the dancer’s pose, (d) remove the extra leg of the dog, and (e) and replace the first dog with artifacts with the second dog.
@@
By supporting the ability to combine generated images, our method allows users to achieve a wider range of results with more flexibility and control. Here, we highlight some use cases and show compelling results for each application. We create the examples using various pre-trained ControlNet models [84] with Stable Diffusion 1.5, such as canny edge, scribble map, Openpose [15], and depth map. Please see Appendix A and E for additional results and details.
Appearance Mixing.   First, we show applications in creative and artistic design, where users refine images based on subjective preference. This is useful in cases where the user may not realize what they want until they see it (e.g., creative exploration). For example, users may use our method for exploring architectural designs, e.g., by combining the roofs, windows, and doors from different images (Figures 5a) or adding features such as a rooftop canon or a glowy courtyard (Figure 5b). Additionally, users can composite different components to create something new. For example, we can combine the body, ear, and arm of a robot to form a new robot (Figure 5e). This strategy can also be applied to other subjects, e.g., to combine new colors in a bird’s feathers (Figure 5d, 15) or a snake’s scales (Figure 5c).
@@
For example, suppose the user wishes to create an Apple-logo-shaped rock and prompts ControlNet with “A rock on grass” alongside an Apple-logo sketch. Since Apple-logo-shaped rocks are not commonly seen in real life (and thus out-of-training distribution), the model fails to produce the desired image. Figure 6a shows an additional rock piece covering the apple bite. To correct it, the user can use our framework to replace it with a patch of grass from another image in the stack (Figure 6a). Similarly, users can correct ControlNet’s output to create waffles in the shape of famous architectural buildings (Figure 6b), which is difficult to achieve with ControlNet alone due to the rare object-shape combination. Finally, we show other correction examples, such as a dancer with an incorrect pose (Figure 6c) and a dog with an extra leg (Figure 6d). Figure 6e shows an example of replacing an unrealistic-looking dog with a different one.
Prompt Alignment.   In addition, our method can be used to increase prompt alignment in cases where the generated output does not accurately follow the input prompt. For example, it is difficult for ControlNet to follow all aspects of long complicated prompts, such as “A red fairy, a green fairy, and a blue fairy sitting from left to right in a brown boat” (Figure 7a). Using our method, users can create the desired image by breaking it up into simpler prompts and selectively combining the outputs (Figure 7b). Since our results depend on the availability of at least one “correct” candidate per region within the generated stack, we encourage our method to be used in conjunction with existing methods [17, 23] for greater accuracy and control.
@@ 5 Evaluation
Our graph cut runs in ${\sim}1$ sec, and the blending stage takes about $7$-$13$ seconds on an NVIDIA A6000 GPU for $2$-$5$ images (of sizes $512\times 512$ to $1024\times 512$). For reference, one forward pass of vanilla ControlNet takes $7$-$17$ sec for a batch of $5$ images. Storage space for the initially generated $QKV$ features depends on image resolution and is ${\sim}2$GB for a $512$x512512512512 image. To save storage space, one could store a subset of features with a trade-off of lower appearance fidelity. Please see Appendix C.2 for more details.
Below, we compare our method against baselines on the examples from Section 4, which spans various use cases.
5.1 Segmentation
Figure 8: Our Graph Cut vs. SAM [43]. We adapt SAM to an image stack as follows: first, we run SAM on each image, using the user strokes as positive and negative samples. Then, for each pixel location, we choose the image label with the largest logit score. As shown, SAM may output noisy, incongruent labels within an object, such as the building (a) and fairies (c). SAM may also fail to follow user strokes, such as the bird interior (b) and the boat (c). Our graph cut takes into account the entire image stack during optimization and outputs labels that are congruent and satisfy user strokes.
@@ 5.1 Segmentation
Figure 9: Qualitative Comparison. Leftmost column: Image regions to blend (output of graph-cut optimization). Graph-cut in diffusion feature space is visualized on the left, and the image-space composite of that graph-cut is visualized on the right. Interactive Digital Photomontage [1]: pixel-space graph-cut may cause seams to fall on undesired edges (see Figure 3 for details), and their gradient-domain blending often fails to preserve color, e.g., the bird’s yellow beak is not preserved in (f). Blended latent diffusion [4] and MasaCtrl+ControlNet [14] may lead to color changes (c, f) and structure changes (a, b, d, e).
Ours Interactive Digital Photomontage Blended Latent Diffusion BLD + MultiDiffusion MasaCtrl + ControlNet Cross-Domain Compositing Deep Image Blending GP-GAN Collage Diffusion
Masked LPIPS $\downarrow$ 0.104 0.085 0.187 0.188 0.198 0.380 0.252 0.220 0.244
PSNR $\uparrow$ 23.44 21.12 21.17 20.66 19.50 20.31 18.35 18.11 20.95
Seam Gradient Score
min: 0.256, avg: 0.337, max: 0.427
0.335 0.312 0.386 0.326 0.341 0.394 0.301 0.207 0.487
@@
Table 1: Quantitative Results. To measure fidelity of local image regions, we report masked LPIPS and PSNR. To measure blending quality of seams, we compute a seam gradient score (SG), which is the mean gradient magnitude along seams that join different image regions. For reference, we also compute the minimum, average, and maximum SG scores of each image stack and report their averages on the left. Methods with SG scores that fall outside of this range exhibit significant seam artifacts (italicized).
We compare our multi-label graph cut with a modified version of SAM [43] as well as the pixel-space graph cut in Interactive Digital Photomontage [1]. Please see Appendix C.1 for an ablation of features used in our graph cut.
SAM.   Because SAM is trained to segment a single image, it is not straightforward to adapt it to a multi-image stack while maintaining coverage and avoiding overlaps. As a baseline, we use the following setup: given an image stack, we run SAM on each image with the user strokes as positive and negative samples. SAM outputs logits per pixel per image. For each pixel, we assign the image label with the largest logit score. As shown in Figure 8, the output labels have two major types of artifacts: 1) incongruent (noisy) labels within an object; 2) the segmentation does not follow user strokes.
Interactive Digital Photomontage.   We compare our feature-based graph cut with their pixel-space counterpart. Specifically, we use their “match edge” seam objective, where the pairwise term $E(p,q,L_{p},L_{q})$ is computed based on a Sobel filter on the RGB values. As shown in Figure 3, the edge strength computed from the filter may be sensitive to low-level changes in color, which could cause their graph cut to select seams that fall on undesired edges. Diffusion feature-based graph cut, on the other hand, selects seams that are more aligned with semantic features. Moreover, as diffusion features are lower resolution than the original images, our graph-cut optimization is more efficient, with $1/64$ the number of variables.
5.2 Blending
For blending, we consider overall realism and local fidelity of user-selected regions. First, we qualitatively compare our results with several baselines:
Pixel-Space Photomontage.   Interactive Digital Photomontage (IDP) uses Poisson blending [59] as a post-process for smoothing composite regions together. Because generated images [59]tend to have wide color variations, gradient-domain blending often leads to changes in color, such as the marked regions in Figure 9. Our method is better at preserving local color and image appearances.
Noise-based Blending.   Blended Latent Diffusion (BLD) [4] blends images by combining their noise at each diffusion step. For this baseline, we inject the noises of image $i=2...N$ into the noise of the base image $i=1$. As shown in Figure 9, BLD may change the appearance of the base image (a, c, e), the injected regions (d), and include artifacts (b, e, f). We also experiment with fusing the noises with greater overlap, inspired by MultiDiffusion [8], but they lead to more artifacts at the seams (see Appendix B).
Attention-based Editing.   We also compare our method with an adapted version of MasaCtrl [14] with ControlNet. We adapt their mask-guided framework by designating the base image as the target image and the other images within the stack as source images. We extend their framework to use multiple foreground masks (one for each image beyond the base image). As shown in Figure 9, this may change the appearance of the base image (a, c, e), selected regions (d, e, f), or show other blending artifacts (b).
Other Blending.   We compare our method against other diffusion-based methods, such as CollageDiffusion [64] and Cross-Domain Compositing [29]. However, these methods fail to preserve local appearance when blending different regions. We also compare with GP-GAN [79] and Deep Image Blending [83], which show artifacts in blended regions. Please see Appendix B for the qualitative results and additional baselines.
@@ 5.2 Blending
Quantitative Metrics.   To measure local appearance fidelity, we computed masked LPIPS [85] and PSNR [33]. For LPIPS, we use the masks from our feature-based graph cut, resized to image dimensions. For PSNR, we compare each blended result with the image-space composite (also created with the graph cut masks). Note that the image-space composites contain noticeable seams (e.g., the first column in Figure 9) and only serve as a proxy in the absence of groundtruth data.
@@
Quantitative Results.   As shown in Table 1, our method achieves the highest PSNR score and the second lowest LPIPS loss across all the baselines. Our method’s SG score is within range and close to the average SG score. CollageDiffusion has a large SG score due to major seam artifacts, and GP-GAN has a low SG score because the images are significantly faded with a reduced color range. Please see Appendix B for additional examples. Deep Image Blending and Cross-Domain Compositing have relatively large LPIPS losses. BLD has the second-highest PSNR score and performs better than its variant, BLD+MultiDiffusion.
@@
Figure 10: User Survey Results. Across $12$ test scenes, we asked participants to select the best image out of four: Interactive Digital Photomontage (DP), Blended Latent Diffusion (BLD), MasaCtrl+ControlNet (MC+CN), and Ours. The results suggest that our method has the best blending quality by a wide margin while being comparable to MasaCtrl+ControlNet in realism.
@@
User Survey.   Finally, we conduct two user surveys to compare our results with the most competitive baseline in each domain: Interactive Digital Photomontage (pixel-space), BLD (noise-based), and MasaCtrl+ControlNet (attention-based). For each test scene, we ask participants to select the best image out of four: (1) Which image appears most realistic to you? and (2) Which image is the best at blending all the selected regions? Results show that our method is comparable to MasaCtrl+ControlNet in terms of realism but has the best blending quality among the baselines by a wide margin (Figure 10).
5.3 Ablation Study
Figure 11: Ablation with Alternative Self-Attention Injection Strategies. (a) Input graph-cut segmentation, visualized here by resizing the diffusion-feature masks to match image dimensions and compositing in pixel space. (b) Inspired by StyleAligned [31], we replace $K^{\text{comp}},V^{\text{comp}}$ with $K^{\text{concat}}=[K^{1},K^{2},...,K^{N}]$ and $V^{\text{concat}}=[V^{1},V^{2},...,V^{N}]$. (c) Similar to Equation 4, we injecte $K^{\text{model}}$ and $V^{\text{model}}$ for the base image. (d) Our method can preserve local image appearances while harmoniously blending them.
Here, we ablate our self-attention injection scheme with alternative injection strategies [31, 2, 14], adapted to our use case. In these comparisons, we use the same graph cut segmentation as input.
First, we consider using shared (concatenated) key and value features, i.e, $K^{\text{concat}}=[K^{1},K^{2},...,K^{N}]$ and $V^{\text{concat}}=[V^{1},V^{2},...,V^{N}]$, which Hertz et al. [31] used to transfer style across different images. Figure 11 shows the results along with the feature-based graph cut, visualized in image space. As shown, the alternative injection schemes can change the local image appearance, such as the dog’s background, the robot’s head, the snake color, and the waffle. Our results align with the observation [2, 14] that $Q$ features influence the image structure, and $K$ and $V$ features influence the image appearance. $K^{\text{concat}},V^{\text{concat}}$ allows $Q$ features to match with keys that are not from the target image region and can thus result in appearance changes. Our method only makes the target image region’s keys and values available and, hence, is able to preserve local appearances.
Next, we consider using $K^{\text{model}}$ and $V^{\text{model}}$ for the base image, similar to Equation 4. Using $K^{\text{model}},V^{\text{model}}$ for the base image also allows appearances to change. By injecting $K^{B}$ and $V^{B}$ of the base image stored during initial generation, our method is able to preserve their appearance during blending. For additional ablations on graph cut features and attention injection parameters, please see Appendix C.
6 Discussion and Limitations

### v3

@@ 1 Introduction
Our framework begins with a stack of images from ControlNet, generated by using the same input condition and different seeds, and lets users choose parts they like from different images via simple brush strokes. Our key insight is that these images share common spatial structures from the same input condition, which can be leveraged for composition. We propose a novel technique that takes in the user’s brush strokes, segments the image parts in diffusion feature space, and then composites these parts during a final denoising process. Specifically, given users’ sparse scribbles, we formulate a multi-label graph-based optimization in diffusion feature space, grouping regions with similar diffusion features while satisfying user inputs. We then introduce a new feature injection and mixing method to composite the segmented regions. Our method accurately preserves the user-selected regions while harmoniously blending them together.
@@
We show visually compelling results on various applications and user workflows, including creating new appearance combinations, correcting shape misalignment, reducing artifacts, and improving prompt alignment. Our method outperforms existing blending methods in preserving the fidelity of local regions while maintaining overall realism. Our code and data are available on our webpage. We include a breadth of results and ablation experiments in the Appendix.
@@ 2 Related Work
Image editing.   Text-to-image diffusion models have enabled various editing tasks given reference images and text-instructions [55, 32, 35, 27, 86, 87, 93]. This is usually achieved through fine-tuning the model [41, 13] or modifying the denoising process of the diffusion model [55, 31, 60, 81, 15, 2, 36]. These methods often focus on the attention mechanism in the text-to-image model, which is crucial for determining the structure and text alignment of generated images [31, 62]. While we take inspiration from these works, our tasks and methods are different. Notably, MasaCtrl [15], Cross-Image Attention [2], and StyleAligned [32] focus on high-level style transfer, where local appearances are expected to change. We focus on blending a multi-image stack and preserving local appearances for greater user control. We show that our proposed technique performs better for this new task.
Controllable image generation.   Improving controllability and adherence to text instructions is critical for using these models as a collaborative tool. As a result, many recent works increase user control in the form of input conditions, such as sketch, depth map, bounding box, segmentation map, and reference image [90, 65, 5, 50, 43, 94, 61, 53, 28, 9]. Another line of work improves the existing text conditioning [18, 24, 7], constrains internal features [80, 6], or augments it through a rich text editor [26]. However, these works aim to create a single correct image by directly constraining the output space of solutions. In contrast, we offer a complementary approach – we allow users to pick and choose exactly what they want from multiple generated images, giving them more fine-grained control without relying solely on the model to create a single perfect image in one shot.
@@ 3.2 Segmentation with Feature-Space Graph Cut
Discussion.   While it is possible to segment each image individually using off-the-shelf methods (e.g., SAM [44]), it often requires extra user intervention to resolve conflicts where objects overlap and to maintain coverage. Instead, our multi-label graph cut optimization can automatically account for all images within the stack, assigning a unique label to each location.
@@ 4 Results
Figure 5: Appearance Mixing and Shape Correction. Our method can be used to mix appearances for creative exploration and design (a, b). Our method can also fix incorrect shapes and artifacts from ControlNet’s outputs (c, d) , which often occur for uncommon input shapes.
@@
By supporting the ability to combine generated images, our method allows users to achieve a wider range of results with more flexibility and control. Here, we highlight some use cases and show compelling results for each application. Due to space constraints, we refer readers to the Appendix for additional results. We created the results using various pre-trained ControlNet models [90] with Stable Diffusion 1.5, such as canny edge, scribble map, Openpose [16], and depth map.
Appearance Mixing.   First, we show applications in creative and artistic design, where users refine images based on subjective preference. This is useful in cases where the user may not realize what they want until they see it (e.g., creative exploration). For example, users may use our method for exploring architectural designs, by combining the roofs, windows, and doors from different images (Figure 1c), or for creating fashion designs, by mixing different features of shoes (Figure 5b). Additionally, users can composite different components to create something new. For example, we can combine the body, ear, and arm of a robot to form a new robot (Figure 1a). This strategy can also be applied to other subjects, e.g., to combine new colors in a bird’s feathers (Figure 5a).
@@
For example, suppose the user wishes to create an Apple-logo-shaped rock and prompts ControlNet with “A rock on grass” alongside an Apple-logo sketch. Since Apple-logo-shaped rocks are not commonly seen in real life (and thus out-of-training distribution), the model fails to produce the desired image. Figure 2 shows an additional rock piece covering the apple bite. To correct it, the user can use our framework to replace it with a patch of grass from another image in the stack (Figure 2b). Similarly, users can correct ControlNet’s output to create waffles in the shape of famous architectural buildings (Figure 5c), which is difficult to achieve with ControlNet alone due to the rare object-shape combination. Finally, we show other correction examples, such as correcting a dancer’s pose (Figure 5d), or replacing an unrealistic-looking dog with a different one (Figure 1d).
Prompt Alignment.   In addition, our method can be used to increase prompt alignment in cases where the generated output does not accurately follow the input prompt. For example, text-to-image with ControlNet can often fail to follow all aspects of a long complicated prompt, such as “A red fairy, a green fairy, and a blue fairy sitting from left to right in a brown boat” (Figure 6a). Using our method, users can create the desired image by breaking it up into simpler prompts and selectively combining the outputs (Figure 6b). Since our results depend on the availability of at least one “correct” candidate per region within the generated stack, we encourage our method to be used in conjunction with existing methods [18, 24] for greater accuracy and control.
@@ 5 Evaluation
For a comprehensive evaluation, we created $20$ test examples, spanning the use cases in Section 4. Below, we compare our method against several baselines. We include the full list of test examples, additional details, and ablations in the Appendix.
@@
Figure 7: Qualitative Comparison. Leftmost column: Image regions to blend (output of graph-cut optimization). Graph-cut in diffusion feature space is visualized on the left, and the image-space composite of that graph-cut is visualized on the right. Interactive Digital Photomontage [1]: its pixel-space graph-cut may cause seams to fall on undesired edges (see Figure 3 also), and their gradient-domain blending often fails to preserve color. Blended Latent Diffusion [4] and MasaCtrl+ControlNet [15] may lead to color and structure changes.
Baselines.   We select baselines that perform blending in pixel space: Interactive Digital Photomontage (IDP) [1]; noise space: Blended Latent Diffusion (BLD) [4]; and attention space: MasaCtrl with ControlNet [15]. We also compare with CollageDiffusion [69], Cross-Domain Compositing (CDC) [30], GP-GAN [85], and Deep Image Blending [89]. In addition, we compare with a modified version of BLD with greater noise overlap, inspired by MultiDiffusion [8]. Please see the Appendix for implementation details. When running comparisons, we use user strokes as input for IDP and our graph cut masks as input for the other baselines.
Masked LPIPS $\downarrow$ Masked SSIM $\uparrow$ PSNR $\uparrow$ Seam Gradient Score min, avg, max 0.255, 0.340, 0.425
Ours 0.123 0.815 22.46 0.339
IDP [1] 0.104 0.888 20.13 0.306
BLD [4] 0.222 0.772 20.27 0.393
BLD+Multi [8] 0.224 0.766 19.79 0.318
MasaCtrl+CtrlNet [15] 0.230 0.680 18.34 0.341
Deep Img Blending [89] 0.270 0.766 17.33 0.313
GP-GAN [85] 0.226 0.820 17.45 0.220*
CDC [30] 0.376 0.584 20.01 0.456*
CollageDiffusion [69] 0.243 0.605 20.57 0.559*
@@
Table 1: Quantitative Results. To measure fidelity of local image regions, we report masked LPIPS and SSIM and PSNR. To measure blending quality of seams, we compute a seam gradient (SG) score, which is the mean gradient magnitude along seams that join different image regions. For reference, we also compute the minimum, average, and maximum SG scores of each image stack and report their averages on the top. Methods with SG scores that fall outside of this range (*) exhibit significant seam artifacts.
@@
Quantitative Metrics.   To measure local appearance fidelity, we computed masked LPIPS [91], masked SSIM [83], and PSNR [34]. For LPIPS and SSIM, we use the masks from our feature-based graph cut, resized to image dimensions. For PSNR, we compare each blended result with the image-space composite (also created with the graph cut masks). Note that the image-space composites contain noticeable seams (e.g., the first column in Figure 7) and only serve as a proxy in the absence of groundtruth data.
@@
Quantitative Results.   As shown in Table 1, our method achieves the highest PSNR score, the second-best LPIPS loss, and the third-best SSIM score across all the baselines, indicating good preservation of local image appearance. While IDP achieved the best LPIPS and SSIM scores, it often exhibits significant color changes (e.g., Figure 7), leading to a lower PSNR score than ours. Our method’s SG score is within range and close to the average SG score. Though CollageDiffusion and GP-GAN achieved the second-best PSNR and SSIM, respectively, they have out-of-range SG scores, denoting major seam artifacts. Deep Image Blending and Cross-Domain Compositing have relatively large LPIPS losses, which correlates with larger changes in image appearances. BLD has the third-highest PSNR score and outperforms its variant, BLD+MultiDiffusion. Please see the Appendix for examples.
@@
Figure 8: User Survey Results. For each test scene, we asked participants to select the best image produced by four methods. The results show that our method has the best blending quality while being comparable to MasaCtrl+CtrlNet in realism. Error bars: SE.
Qualitative Comparison.   In Figure 7, we show qualitative comparisons with IDP, BLD, and MasaCtrl+ControlNet. As shown, IDP often leads to changes in color due to gradient-domain blending, such as the marked regions in Figure 7. BLD [4] and MasaCtrl+ControlNet [15] may lead to structural changes or local appearance changes. The other baselines exhibit significant artifacts in blended regions or at seams. Please see the Appendix for qualitative results of these other baselines.
@@
User Survey.   Finally, we conduct two user surveys to compare our results with the most competitive baseline in each domain: Interactive Digital Photomontage (pixel-space), BLD (noise-based), and MasaCtrl+ControlNet (attention-based). Across $12$ test scenes, we ask participants to select the best image produced by the four methods: (1) Which image appears most realistic to you? and (2) Which image is the best at blending all the selected regions? Out of $324$ and $240$ responses, respectively, results show that our method is comparable to MasaCtrl+ControlNet in terms of realism and has the best blending quality among the baselines by a wide margin (Figure 8).
5.1 Ablation
Masked LPIPS $\downarrow$ Masked SSIM $\uparrow$ PSNR $\uparrow$ Seam Gradient Score min, avg, max 0.255, 0.340, 0.425
Ours 0.123 0.815 22.46 0.339
w/ $K^{\text{concat}}$, $V^{\text{concat}}$ 0.243 0.677 18.37 0.354
w/ $K^{\text{model}}$, $V^{\text{model}}$ 0.268 0.669 18.85 0.332
Table 2: Ablation of Self-Attention Injection Schemes.
Here, we ablate our self-attention injection scheme with alternative injection strategies, adapted to our use case. First, we consider using shared (concatenated) key and value features, i.e, $K^{\text{concat}}=[K^{1},K^{2},...,K^{N}]$ and $V^{\text{concat}}=[V^{1},V^{2},...,V^{N}]$, which StyleAligned [32] used to transfer style across different images. Next, we consider using $K^{\text{model}}$ and $V^{\text{model}}$ for the base image, similar to Equation 4. Both alternative injection schemes lead to changes in the appearance of user-selected image regions, whereas our method is able to preserve them. Table 2 reflects this change quantitatively. Please see the Appendix for qualitative examples and additional ablations on attention injection.
5.2 Multi-image Segmentation
Figure 9: Our Graph Cut vs. SAM [44]. We adapted SAM to an image stack and compared it with our graph cut. As shown, SAM may output noisy, incongruent labels within an object, such as the building (a) and fairies (c). SAM may also fail to follow user strokes, such as the bird interior (b) and the boat (c). Our graph cut takes into account the entire image stack during optimization and outputs labels that are congruent and satisfy user strokes.
Here, we compare our multi-label graph cut with a modified version of SAM [44] and pixel-space graph cut in IDP [1]. Since SAM is trained to segment single images, adapting it to a multi-image stack while maintaining coverage and avoiding overlaps is not straightforward. As a baseline, we use the following setup: given an input image stack, we run SAM on each image, with the user strokes on the image as positive and strokes on other images as negative samples. SAM outputs logits per pixel per image. For each pixel in the composite image, we assign the image label with the highest logit score. As shown in Figure 9, we observe two major types of artifacts here: 1) incongruent (noisy) labels within an object; 2) the segmentation does not follow user strokes. Pixel-space graph cut, as shown in Figure 3, may be sensitive to low-level changes in color and select seams that fall on undesired edges. Our graph cut, which operates in feature space, selects seams that are more aligned with semantic features. Please see the Appendix for ablations of these methods and other graph cut features.
6 Discussion